# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

### v2

@@ 1 Introduction
Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information gpt4_technical_report ; gpt3_technical_report ; llama3_technical_report . Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand. For example, OpenAI’s production-scale models process roughly 100 billion words per day openai_billion_token and production query rates frequently surpass 200 requests per second chatgpt_fact ; scale_llm .
To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput orca ; paged_attention but also on meeting service-level objectives (SLOs) sarathi_serve ; llumnix ; sola_llm . One of the key challenges in achieving SLOs for LLM serving lies in handling dynamic load. Request rates, as well as input and output sequence lengths, can fluctuate significantly over short time intervals. Simply overprovisioning compute resources may mitigate this issue but results in substantial system cost inefficiency.
As a promising alternative, we explore an approach we call dual-precision LLM serving, which utilizes FP8 together with FP16, the de facto standard number format for weight parameters. While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 . By operating in FP16 mode under normal conditions and switching to FP8 mode only during sudden load surges, we can effectively manage dynamic workloads without resource overprovisioning. This approach achieves a better balance between SLO attainment and service quality than relying solely on either FP16 or FP8.
@@
In this paper, we present NestedFP, a technique that enables efficient dynamic selection between FP8 and FP16 precision from a single unified model representation. NestedFP decomposes each 16-bit weight tensor into two 8-bit components, allowing FP8 inference without additional memory overhead or throughput degradation. Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library cutlass , that performs accurate on-the-fly FP16 reconstruction during kernel execution. According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.47% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
@@ 2.1 Floating Point Representation
Commonly used 16-bit formats include FP16 (E5M10) and BF16 (E8M7), while prevalent 8-bit formats include E4M3 and E5M2. These formats are natively supported by many modern hardware platforms and have become standard choices in low-precision machine learning computations investigation_fp8_across_accelerators .
@@ 2.2 FP8 Quantization for LLMs
First, floating-point quantization is generally better suited for LLMs than integer quantization due to its wider dynamic range. Unlike in computer vision tasks—where integer quantization is prevalent—LLMs often exhibit activation outliers amplified by operations such as LayerNorm efficient_ptq_fp8 . These outliers can significantly degrade model performance if not properly managed investigation_fp8_across_accelerators . Although recent work has attempted to mitigate this issue in integer quantization through static outlier-aware techniques efficient_ptq_fp8 ; smoothquant , these methods still underperform compared to floating-point quantization, which inherently better accommodates long-tailed distributions and high dynamic ranges llm_fp4 ; fp8_formats_for_deep_learning ; int_or_fp .
@@
Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators . Leveraging this support, FP8 GEMM operations can achieve up to $2\times$ speedup compared to their FP16 counterparts. This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
@@
E4M3 Format. Among the two widely adopted FP8 formats—E4M3 and E5M2—existing studies have consistently shown that E4M3 yields higher inference accuracy for Transformer-based language models fp8_formats_for_deep_learning ; efficient_ptq_fp8 ; hybrid_fp8_training_and_inference . However, due to its reduced dynamic range compared to FP16, the effectiveness of E4M3 critically depends on the application of appropriate scaling strategies the_power_of_exponent ; fp8_formats_for_deep_learning ; hybrid_fp8_training_and_inference . Scaling is applied to align the dynamic range of tensor values with the numerical bounds of the E4M3 format. Scaling factors may be determined statically—based on heuristics such as the absolute maximum, percentile clipping, or mean squared error (MSE) minimization fp8_formats_for_deep_learning —or computed dynamically during runtime. Weight tensors are typically scaled statically on a per-channel basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp , most commonly using the absolute maximum value. Activation tensors, by contrast, may be scaled offline on a per-tensor basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp or dynamically on a per-token basis during inference. While the latter approach can improve accuracy, it incurs additional runtime cost due to the overhead of computing scaling factors on-the-fly investigation_fp8_across_accelerators ; training_and_inference_llm_fp8 .
@@
FP8 Results. While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision dfloat11 . Table 1 presents a comparison between FP8 and FP16 across various models and benchmarks. Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm . A consistent accuracy degradation is observed in FP8 models, which clearly shows the trade-off between throughput and model quality in FP8 quantization.
@@ 3.1 Challenge: Dynamic Load in LLM Serving
LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen . These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in Time To First Token (TTFT) and Time Per Output Token (TPOT). Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet SLOs.
To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace . Specifically, we measure how the request rate fluctuates over a one-hour interval (00:00–01:00 UTC, May 10, 2024, in the trace). Figure 1 presents the results. The request rate exhibits nearly a fivefold variation between the lowest and highest load periods. Moreover, as shown in the one-minute zoom-in view, these fluctuations occur at high temporal frequency, indicating second-level variability.
A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) auto_scaling . However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales auto_scaling_timescale . Consequently, it is inadequate for LLM serving systems, which must respond to sharp, second-level load spikes and scale throughput within millisecond-level intervals to meet SLO requirements.
@@ 3.2 Our Proposal: Dual-Precision LLM Serving
Figure 1 presents the per-second p90 TPOT under three precision schemes: FP16, FP8, and the proposed dual-precision format, using the Llama 3.1 8B model on an H100 GPU with vLLM. The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace azure_llm_inference_trace , scaled down to 20% of the original load. The TPOT SLO threshold is set to 33 ms baseten_slo . As shown, the FP16 model exhibits more frequent TPOT spikes that exceed the TPOT SLO threshold compared to the FP8 model. To further quantify the difference, we analyze the rate of SLO violations across the same trace: 35.8% of requests under FP16 exceeded the threshold, whereas FP8 reduced this to 17.3%. In short, the proposed dual-precision inference can achieve FP8-level SLO compliance while minimizing quality degradation.
@@ 4.1 Compact yet Effective Data Format for Dual-Precision LLM
Handling Exception Layers. For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is non-zero), we retain them in FP16 and always perform FP16 computation. This may reduce the benefits of dual-precision LLM serving, as the throughput advantage of FP8 mode becomes less pronounced. However, it is important to recall that, as shown in our analysis of popular models (Figure 3), such layers are rare and therefore have minimal impact on the overall effectiveness of the dual-precision scheme.
@@ 4.2 FP16 GEMM Kernel with On-the-fly Reconstruction
Efficient FP16 Reconstruction via Merged Bitwise Operations. FP16 reconstruction involves a large number of 8-bit bitwise operations. To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in any_precision_llm ; fp6_llm . Algorithm 1 illustrates how FP16 reconstruction is performed using merged bitwise operations. Conceptually, the process iterates over the weight tensor, where each iteration processes four elements. In each iteration, four 8-bit values are loaded from the upper and lower tensors ($W_{\text{upper}}$ and $W_{\text{lower}}$) into 32-bit registers u and l, respectively (Line 5–6). Then, the sign bits are extracted from u and stored in another register s (Line 7). Next, the algorithm checks whether rounding has occurred by examining l (Line 8); if rounding is detected, the effect is reversed, thereby recovering the exponent and mantissa parts of the upper 8 bits (Line 9). Then, the original upper 8 bits, u_orig, are fully reconstructed by concatenating them with s (Line 10). Finally, u_orig and l are fused using NVIDIA’s __byte_perm instruction (lines 11–12), completing the reconstruction. Note that all the above operations are simultaneously applied to four 8-bit values through a 32-bit instruction, effectively reducing the total number of bitwise operations required.
@@ 5.1 FP8 Model Quality
Methodology. We evaluate the accuracy of the FP8 model in NestedFP against FP8 models quantized using commonly adopted configurations, specifically per-channel weight quantization. For activation quantization, we apply per-token quantization for both NestedFP and the baseline. The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B deepseek_ref ; nemo_ref ; small_ref ; llama3_technical_report ; phi_ref . For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH minerva ; bbh ; mmlupro . All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness lm-evaluation-harness . As noted in Section 4.1, 8.75% of the linear layers in Phi-4 are not applicable to NestedFP and are therefore executed in FP16. More details on the experimental setup are provided in Appendix A.
@@
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small Llama 3.1 70B DeepSeek-R1- Distill-Llama-70B FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.5 16.4 42.9 42.9 34.2 35.2 35.7 34.6 46.0 45.5 MMLU Pro 32.8 33.2 35.3 34.7 52.7 53.3 53.6 53.7 48.1 47.4 49.2 49.4 BBH 38.6 39.0 40.6 40.0 27.1 28.7 50.8 51.1 49.1 49.7 53.2 53.4
Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks, compared to the baseline FP8 model. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
@@ 5.3 FP16 Mode End-to-End Inference Performance
Methodology. We evaluate the end-to-end inference throughput of the FP16 mode of NestedFP by integrating it into the vLLM framework on a single NVIDIA H100 GPU. Specifically, we set the input and output sequence lengths to 1024 and 512, respectively, while varying the batch size from 32 to 512. More details on the experimental setup are provided in Appendix C. The results are compared against vanilla vLLM FP16 execution, which relies on PyTorch for invoking GEMM kernels. Again, the FP8-mode throughput of NestedFP is identical to baselines such as vanilla vLLM and is therefore omitted from evaluation.
Results. Figure 7 presents the results. NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (LLaMA 3.1–8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes. The end-to-end throughput degradation is even lower than the kernel-level overheads reported in Section 5.2, since non-GEMM components also contribute to total execution time. Additional end-to-end throughput results are provided in Appendix C.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information Achiam et al. (2023); Brown et al. (2020); Meta (2024). Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand. For example, OpenAI’s production-scale models process roughly 100 billion words per day Griffin (2024) and production query rates frequently surpass 200 requests per second Lammertyn ; Yao et al. (2024).
To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput Kwon et al. (2023); Yu et al. (2022) but also on meeting service-level objectives (SLOs) Agrawal et al. (2024); Hong et al. (2025); Sun et al. (2024). One of the key challenges in achieving SLOs for LLM serving lies in handling dynamic load. Request rates, as well as input and output sequence lengths, can fluctuate significantly over short time intervals. Simply overprovisioning compute resources may mitigate this issue but results in substantial system cost inefficiency.
As a promising alternative, we explore an approach we call dual-precision LLM serving, which utilizes FP8 together with FP16, the de facto standard number format for weight parameters. While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 Micikevicius et al. (2022); Perez et al. (2023); Shen et al. (2024). By operating in FP16 mode under normal conditions and switching to FP8 mode only during sudden load surges, we can effectively manage dynamic workloads without resource overprovisioning. This approach achieves a better balance between SLO attainment and service quality than relying solely on either FP16 or FP8.
@@
In this paper, we present NestedFP, a technique that enables efficient dynamic selection between FP8 and FP16 precision from a single unified model representation. NestedFP decomposes each 16-bit weight tensor into two 8-bit components, allowing FP8 inference without additional memory overhead or throughput degradation. Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library NVIDIA , that performs accurate on-the-fly FP16 reconstruction during kernel execution. According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.38% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
@@ 2.1 Floating Point Representation
Commonly used 16-bit formats include FP16 (E5M10) and BF16 (E8M7), while prevalent 8-bit formats include E4M3 and E5M2. These formats are natively supported by many modern hardware platforms and have become standard choices in low-precision machine learning computations Kim et al. (2025).
@@ 2.2 FP8 Quantization for LLMs
First, floating-point quantization is generally better suited for LLMs than integer quantization due to its wider dynamic range. Unlike in computer vision tasks—where integer quantization is prevalent—LLMs often exhibit activation outliers amplified by operations such as LayerNorm Shen et al. (2024). These outliers can significantly degrade model performance if not properly managed Kim et al. (2025). Although recent work has attempted to mitigate this issue in integer quantization through static outlier-aware techniques Shen et al. (2024); Xiao et al. (2023), these methods still underperform compared to floating-point quantization, which inherently better accommodates long-tailed distributions and high dynamic ranges Liu et al. (2023); Micikevicius et al. (2022); Zhang et al. (2024).
@@
Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs Kim et al. (2025). Leveraging this support, FP8 GEMM operations can achieve up to $2\times$ speedup compared to their FP16 counterparts. This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
@@
E4M3 Format. Among the two widely adopted FP8 formats—E4M3 and E5M2—existing studies have consistently shown that E4M3 yields higher inference accuracy for Transformer-based language models Micikevicius et al. (2022); Shen et al. (2024); Sun et al. (2019). However, due to its reduced dynamic range compared to FP16, the effectiveness of E4M3 critically depends on the application of appropriate scaling strategies Kuzmin et al. (2022); Micikevicius et al. (2022); Sun et al. (2019). Scaling is applied to align the dynamic range of tensor values with the numerical bounds of the E4M3 format. Scaling factors may be determined statically—based on heuristics such as the absolute maximum, percentile clipping, or mean squared error (MSE) minimization Micikevicius et al. (2022)—or computed dynamically during runtime. Weight tensors are typically scaled statically on a per-channel basis Liu et al. (2023); Shen et al. (2024); Zhang et al. (2024), most commonly using the absolute maximum value. Activation tensors, by contrast, may be scaled offline on a per-tensor basis Liu et al. (2023); Shen et al. (2024); Zhang et al. (2024) or dynamically on a per-token basis during inference. While the latter approach can improve accuracy, it incurs additional runtime cost due to the overhead of computing scaling factors on-the-fly Kim et al. (2025); Perez et al. (2023).
@@
FP8 Results. While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision Zhang et al. . Table 1 presents a comparison between FP8 and FP16 across various models and benchmarks. Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm-project . A consistent accuracy degradation is observed in FP8 models, which clearly shows the trade-off between throughput and model quality in FP8 quantization.
@@ 3.1 Challenge: Dynamic Load in LLM Serving
LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests Liu et al. (2024); Sun et al. (2025). These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in Time To First Token (TTFT) and Time Per Output Token (TPOT). Such fluctuations occur in a highly dynamic and often unpredictable manner Sun et al. (2024), making it difficult for service providers to consistently meet SLOs.
To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service Azure (2024). Specifically, we measure how the request rate fluctuates over a one-hour interval (00:00–01:00 UTC, May 10, 2024, in the trace). Figure 1 presents the results. The request rate exhibits nearly a fivefold variation between the lowest and highest load periods. Moreover, as shown in the one-minute zoom-in view, these fluctuations occur at high temporal frequency, indicating second-level variability.
A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) Catillo et al. (2023). However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales Dogani et al. (2023). Consequently, it is inadequate for LLM serving systems, which must respond to sharp, second-level load spikes and scale throughput within millisecond-level intervals to meet SLO requirements.
@@ 3.2 Our Proposal: Dual-Precision LLM Serving
Figure 1 presents the per-second p90 TPOT under three precision schemes: FP16, FP8, and the proposed dual-precision format, using the Llama 3.1 8B model on an H100 GPU with vLLM. The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace Azure (2024), scaled down to 20% of the original load. The TPOT SLO threshold is set to 33 ms Baseten (2025). As shown, the FP16 model exhibits more frequent TPOT spikes that exceed the TPOT SLO threshold compared to the FP8 model. To further quantify the difference, we analyze the rate of SLO violations across the same trace: 35.8% of requests under FP16 exceeded the threshold, whereas FP8 reduced this to 17.3%. In short, the proposed dual-precision inference can achieve FP8-level SLO compliance while minimizing quality degradation.
@@ 4.1 Compact yet Effective Data Format for Dual-Precision LLM
Handling Exception Layers. For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is possibly non-zero), we retain them in FP16 and always perform FP16 computation. This may reduce the benefits of dual-precision LLM serving, as the throughput advantage of FP8 mode becomes less pronounced. However, it is important to recall that, as shown in our analysis of popular models (Figure 3), such layers are rare and therefore have minimal impact on the overall effectiveness of the dual-precision scheme.
@@ 4.2 FP16 GEMM Kernel with On-the-fly Reconstruction
Efficient FP16 Reconstruction via Merged Bitwise Operations. FP16 reconstruction involves a large number of 8-bit bitwise operations. To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in Park et al. (2024); Xia et al. (2024). Algorithm 1 illustrates how FP16 reconstruction is performed using merged bitwise operations. Conceptually, the process iterates over the weight tensor, where each iteration processes four elements. In each iteration, four 8-bit values are loaded from the upper and lower tensors ($W_{\text{upper}}$ and $W_{\text{lower}}$) into 32-bit registers u and l, respectively (Line 5–6). Then, the sign bits are extracted from u and stored in another register s (Line 7). Next, the algorithm checks whether rounding has occurred by examining l (Line 8); if rounding is detected, the effect is reversed, thereby recovering the exponent and mantissa parts of the upper 8 bits (Line 9). Then, the original upper 8 bits, u_orig, are fully reconstructed by concatenating them with s (Line 10). Finally, u_orig and l are fused using NVIDIA’s __byte_perm instruction (lines 11–12), completing the reconstruction. Note that all the above operations are simultaneously applied to four 8-bit values through a 32-bit instruction, effectively reducing the total number of bitwise operations required.
@@ 5.1 FP8 Model Quality
Methodology. We evaluate the accuracy of the FP8 model in NestedFP against FP8 models quantized using commonly adopted configurations, specifically per-channel weight quantization. For activation quantization, we apply per-token quantization for both NestedFP and the baseline. The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B AI (2024, 2025); DeepSeek-AI ; Meta (2024); Microsoft . For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH Lewkowycz et al. (2022); Suzgun et al. (2023); Wang et al. (2024). All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness EleutherAI . As noted in Section 4.1, 8.75% of the linear layers in Phi-4 are not applicable to NestedFP and are therefore executed in FP16. More details on the experimental setup are provided in Appendix A.
@@
Minerva Math MMLU Pro BBH
Model Size FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N)
Llama 3.1 8B 17.6 17.2 32.8 33.2 38.6 39.0
Mistral Nemo 12B 16.5 16.4 35.3 34.7 40.6 40.0
Phi-4 14B 42.9 42.9 52.7 53.3 27.1 28.7
Mistral Small 24B 34.2 35.2 53.6 53.7 50.8 51.1
Llama 3.1 70B 35.7 34.6 48.1 47.4 49.1 49.7
DeepSeek-R1-Distill-Llama 70B 46.0 45.5 49.2 49.4 53.2 53.4
Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
@@ 5.3 FP16 Mode End-to-End Inference Performance
Methodology. We evaluate the end-to-end inference throughput of the FP16 mode of NestedFP by integrating it into the vLLM framework on a single NVIDIA H100 GPU. Specifically, we set the input and output sequence lengths to 1024 and 512, respectively, while varying the batch size from 32 to 512. More details on the experimental setup are provided in Appendix C. The results are compared against vanilla vLLM FP16 execution, which relies on PyTorch for invoking GEMM kernels. The FP8-mode throughput of NestedFP is excluded from the evaluation, as it does not require a separate kernel.
Results. Figure 7 presents the results. NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (Llama 3.1 8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes. The end-to-end throughput degradation is even lower than the kernel-level overheads reported in Section 5.2, since non-GEMM components also contribute to total execution time. Additional end-to-end throughput results are provided in Appendix C.